An iron cube has the volume of 4.76 cm^3. If iron has a density of 7.680 x 10^3 kg/m^3, how many moles are in the cube?


1
Expert's answer
2021-02-23T04:36:45-0500

We have given,


Volume of Iron Cube = "4.76cm^{3}" = "4.76\times10^{-3}m^{3}"


Density = "7.680\times10^{3}kg/m^{3}"


We know,


Density ="\dfrac{Mass}{Volume}"


"7.680\times10^{3} = \dfrac{Mass}{4.76\times10^{-3}}"


Mass = "36.55kg"


Hence, number of moles in the cube= "\dfrac{Mass}{Molar Mass}"


= "\dfrac{36.55}{55.84}"


= "0.6545 moles"